# DK-SLAM: Monocular Visual SLAM with Deep Keypoints Adaptive Learning, Tracking and Loop-Closing

### v1

@@ Abstract
Unreliable feature extraction and matching in handcrafted features undermine the performance of visual SLAM in complex real-world scenarios. While learned local features, leveraging CNNs, demonstrate proficiency in capturing high-level information and excel in matching benchmarks, they encounter challenges in continuous motion scenes, resulting in poor generalization and impacting loop detection accuracy. To address these issues, we present DK-SLAM, a monocular visual SLAM system with adaptive deep local features. MAML optimizes the training of these features, and we introduce a coarse-to-fine feature tracking approach. Initially, a direct method approximates the relative pose between consecutive frames, followed by a feature matching method for refined pose estimation. To counter cumulative positioning errors, a novel online learning binary feature-based online loop closure module identifies loop nodes within a sequence. Experimental results underscore DK-SLAM’s efficacy, outperforms representative SLAM solutions, such as ORB-SLAM3 on publicly available datasets.
@@ I Introduction
Visual localization and mapping constitute crucial components within intelligent mobile agents for motion tracking and environment understanding, with diverse applications ranging from autonomous vehicles, unmanned aerial vehicles (UAVs), and robots to immersive technologies such as Virtual Reality (VR) and Augmented Reality (AR). The imperative nature of achieving robust localization for unmanned platforms across various environmental conditions has garnered escalating attention within the research community.
A Visual Simultaneous Localization and Mapping (SLAM) system comprises front-end perception and back-end optimization. The front-end uses sensors for environmental perception, employing keypoint features in stages like extraction, matching, outlier removal, and pose estimation. The back-end includes loop detection and global map optimization, refining the system state with geometric models. Visual SLAM methods are categorized as keypoint-based and direct methods. Direct methods optimize pixel-level photometric loss but can be unstable due to lighting variations [1, 2, 3]. Keypoint-based methods, like ORB-SLAM2 and ORB-SLAM3 [4, 5], enhance stability by using detectors and descriptors, although handcrafted descriptors are susceptible to lighting and texture variations. Addressing these challenges is crucial for improving the reliability of visual localization and mapping systems in diverse applications.
In recent years, deep learning has shown impressive performance in image processing, particularly in image classification and object detection. The application of deep learning to VO/SLAM methods has gained attention, with two predominant approaches: end-to-end and learned features based methods. End-to-end methods utilize Deep Neural Networks (DNNs) to directly convert images into pose estimates [6]. However, they exhibit low interpretability and suboptimal generalization. In contrast, learned-feature-based SLAM replaces local features, allowing deep learning to focus on feature perception [7, 8, 9, 10]. This approach combines the strengths of deep learning and geometric models but may still face challenges in generalization. Additionally, it may lack crucial low-level keypoint information, impacting feature matching performance. The training complexity of learned-feature-based Bag-of-Words (BoW) adds further challenges, often requiring offline training and exacerbating generalization issues.
To address the aforementioned challenges, we introduce DK-SLAM, a novel Deep Keypoints based SLAM system incorporating adaptive learning, coarse-to-fine feature tracking, and online loop-closing capabilities. This monocular visual SLAM framework utilizes a modified MAML-based training strategy for robust deep feature extraction. It enhances feature matching precision by incorporating patch pixels and constructs a photometric loss using adjacent frame patches to optimize coarse relative pose estimates. An online BoW model with binary deep features adapts dynamically to new environments, mitigating generalization challenges in deep learning, and improves adaptability without requiring additional training. Our proposed DK-SLAM outperforms the representative baselines, e.g. ORB-SLAM3, LDSO and LIFT-SLAM, on the public KITTI and EuRoC datasets.
@@
  • We propose DK-SLAM, a monocular SLAM system with adaptive deep keypoints learning. Our deep feature extractor, trained using a MAML strategy, enhances adaptability to diverse scenes.
  • Our method employs a coarse-to-fine feature matching strategy, estimating relative poses through patch photometric loss optimization and refining them based on the 2D-3D relationship for improved accuracy.
  • We introduce an deep keypoints online learning-based Bag-of-Words (BoW) models that overcomes poor generalization, ensuring correct loop detection in the sequence.
@@ II Related Works
II-A Deep Feature Extractor for Visual SLAM
@@
Mainstream handcrafted extractors, including ORB [11], SIFT [12], and Shi-Thomas [13], are employed in visual SLAMs like ORB-SLAM2 [4] and VINS-Mono [14]. However, these methods relying on gradient information face challenges in dynamic lighting and low-texture environments. To address these issues, researchers explored deep feature extractors. Superpoint [15] employs a shared encoder for high-level information and distinct decoders for keypoints and descriptors, using synthesized labels for self-supervised training. D2-Net [16] eliminates the need for a separate keypoint decoder, ensuring robustness. To enhance repeatability detection, [17] combines multi-scale gradients with depth feature maps. Additionally, L2-Net [18] introduces a local matching loss based on the L2 distance, improving feature matching performance. Experimental results demonstrate the effective feature matching performance of the learned local features on public datasets. Despite their success, the mentioned deep extractors exhibit inherent flaws in the face of diverse scenes, leading to potential performance degradation.
II-B Hybrid methods for Visual SLAM
@@
Recent research explores hybrid SLAM methods to combine deep learning and multi-view geometry advantages. GCNv2-SLAM integrates GCN [19] into SLAM, constructing a complete system with offline BoW training. In [20], knowledge distillation from HF-Net [21] enhances feature detection in a compact model. [22] uses a self-supervised thermal optical flow tracking network, while [10] builds on Superpoint [15] with edge and multi-scale information. Our proposed SLAM methodology shares similarities with [8]. In [8], a LIFT [23] extractor serves as the SLAM front-end, implementing an adaptive feature matching strategy tailored to different datasets. In contrast, our SLAM adopts a meta-learning based training strategy, eliminating the need for separate hyperparameter designs for distinct datasets.
II-C Loop Closure Module
Loop closure is a critical module of visual SLAM, responsible for detecting loop nodes and refining initial pose estimates. The BoW model, such as DBow2, DBow3, and FBow [24], relies on the binary descriptor BRIEF [25]. DBow utilizes a k-d tree structure based on Hamming distance for descriptor storage. However, conventional BoW methods require extensive offline training data, prompting the introduction of iBow [26]. iBow achieves online word bag training using testing datasets, demonstrating robust transfer performance across various scenarios and addressing adaptability concerns. Another closed-loop detection method, akin to NetVLAD [27] in [20, 9], is highly reliant on the training set. Substantial differences between test and training sets may result in performance degradation.
@@ III Deep Keypoint based Monocular SLAM
An overview of our proposed DK-SLAM is depicted in Figure 2. We employ an adaptive local feature extractor, based on Superpoint, for extracting keypoints across multi-scale image pyramids. To balance the number of keypoints on varying scale images, we implement an averaging distribution strategy. For enhanced generalization, the local feature extractor undergoes MAML during training. Subsequently, we optimize memory consumption by converting floating descriptors into binary hash codes. Lastly, an online BoW module is introduced to establish a loop closure mechanism, facilitating closed-loop detection and global map optimization.
Figure 1: An overview of our proposed DK-SLAM framework with adaptive keypoints, two-stage tracking and loop closure.
III-A Adaptive Local Feature Learning
@@ III-A1 Feature Extractor Network
Inspired by Superpoint, our deep local feature extractor also utilizes VGG16 [28] as the backbone. Different from Superpoint, we introduce BatchNorm layers after each Convolutional Neural Network (CNN) to enhance training convergence. Post the backbone, two heads emerge, each comprising multi-layer CNNs. The first, the detector head, generates a probability mask matching the input image size. We posit that positions with a probability surpassing a set threshold on the mask denote keypoint positions. The second head, the descriptor head, predicts feature descriptors for keypoints, normalized to a dimension of 256 using the L2 norm.
III-A2 Self-supervised Adaptive Feature Learning
Our network undergoes self-supervised training, employing a multi-step process. Initially, we obtain the homography-warped image from current training image. To train keypoint detector, we leverage the pre-trained MagicPoint network [15] to acquire keypoint pseudo labels for both the original and warped images. In both images, keypoint predictions and pseudo labels contribute to the formation of detector loss, denoted as ${L_{p}}$ and ${L_{wp}}$, with CrossEntropy serving as the metric. Moving on to descriptor training, we utilize the sparse descriptor loss ${L_{d}}$ proposed in [29]. This involves obtaining both the warped image and pixel-wise correspondences. Descriptor loss functions are then constructed solely for pseudo-labeled keypoints, employing triplet loss. The sparse descriptor loss is composed of two components: homography-matched descriptor loss ($L_{dm}$) and non-matched descriptor loss ($L_{dn}$). The homography-matched descriptor loss, represented by Equation 1, involves a subscript $i$ ranging between $(0,1,...,N)$, where $(\mathbf{p}_{i},\hat{\mathbf{p}_{i}})$ denotes the $N$ matched keypoint position pairs.
$L_{dm}(\mathbf{p}_{i})={\left\|\mathbf{d}_{I}(\mathbf{p}_{i})-\mathbf{d}_{wI}(\hat{\mathbf{p_{i}}})\right\|}^{2}$ (1)
As depicted in Equation 2, we use $M$ pairs of descriptors from the vicinity of matched descriptors to formulate a non-matched descriptor loss. The subscript $j$ ranges from 0 to $M$, where $(\mathbf{p}_{i},\mathbf{p^{\prime}}_{j})$ represents unmatched keypoint position pairs.
$L_{dn}(\mathbf{p}_{i})=\frac{1}{M}\sum_{j=1}^{M}(\left\|\mathbf{d}_{I}\mathbf{(p_{i})}-\mathbf{d}_{wI}(\mathbf{p^{\prime}}_{j})\right\|^{2})$ (2)
@@ III-A2 Self-supervised Adaptive Feature Learning
III-A3 MAML for Feature Extractor Training
@@
Learning-based local feature extractors face poor generalization due to variances in visual features across scenes. Significant feature disparity between training and unseen datasets can lead to catastrophic forgetting. To address this, we leverage insights from Model-Agnostic Meta-Learning (MAML), adapting network parameters through meta-training for improved generalization and rapid adaptation to new data while preventing catastrophic forgetting. The MAML-based meta-training consists of inner loop training and outer loop training. The training set is partitioned into a support set $\mathbf{D}_{s}$ and a query set $\mathbf{D}_{q}$, where $D_{s}$ is involved in inner loop training, and $\mathbf{D}_{q}$ participates in outer loop training.
@@
In the MAML training strategy, the original network parameter $\mathbf{\theta}_{a}$ is initially duplicated to $\mathbf{\theta}_{b}$. Subsequently, both parameters undergo training in both the inner and outer loops. Within the inner loop, the support dataset $\mathbf{D}_{s}$ is utilized for iterative updates to $\mathbf{\theta}_{b}$. This involves dividing a batch of support dataset into $n$ distinct tasks $\tau$, performing $m$ parameter updates on each task. Consequently, within one inner loop, the parameters are updated $n*m$ times, resulting in the updated parameter $\mathbf{\theta}^{n*m}_{b}$. After completing one iteration of inner loop training, a complete batch of query set data is used to train the network. Although the gradient $\mathbf{\bigtriangledown_{\theta^{n*m}_{b}}}$ is calculated for $\mathbf{\theta}^{n*m}_{b}$ in the query set data, it is not used to update $\mathbf{\theta}^{n*m}_{b}$; instead, the original $\mathbf{\theta}_{a}$ is updated.
III-A4 Keypoint Distribution Strategy

### v2

# DK-SLAM: Monocular Visual SLAM with Deep Keypoint Learning, Tracking and Loop-Closing

@@ Abstract
The performance of visual SLAM in complex, real-world scenarios is often compromised by unreliable feature extraction and matching when using handcrafted features. Although deep learning-based local features excel at capturing high-level information and perform well on matching benchmarks, they struggle with generalization in continuous motion scenes, adversely affecting loop detection accuracy. Our system employs a Model-Agnostic Meta-Learning (MAML) strategy to optimize the training of keypoint extraction networks, enhancing their adaptability to diverse environments. Additionally, we introduce a coarse-to-fine feature tracking mechanism for learned keypoints. It begins with a direct method to approximate the relative pose between consecutive frames, followed by a feature matching method for refined pose estimation. To mitigate cumulative positioning errors, DK-SLAM incorporates a novel online learning module that utilizes binary features for loop closure detection. This module dynamically identifies loop nodes within a sequence, ensuring accurate and efficient localization. Experimental evaluations on publicly available datasets demonstrate that DK-SLAM outperforms leading traditional and learning based SLAM systems, such as ORB-SLAM3 and LIFT-SLAM. These results underscore the efficacy and robustness of our DK-SLAM in varied and challenging real-world environments.
@@ I Introduction
Visual localization and mapping are fundamental components in autonomous systems for motion estimation and environmental perception. These capabilities have diverse applications, spanning from self-driving vehicles and unmanned aerial vehicles (UAVs) to mobile robots and immersive wearable technologies such as virtual reality (VR) and augmented reality (AR) devices. The quest for robust localization and mapping under varying environmental conditions has increasingly drawn the research community’s attention.
Visual Simultaneous Localization and Mapping (SLAM) systems generally comprise two main components: front-end perception and back-end optimization. They are typically divided into two categories: direct methods and feature-based methods. Direct methods, such as LSD-SLAM [1], DSO [2], and SVO [3], process the pixel intensities in an image to optimize a photometric loss at the pixel level, directly using image data for map construction and pose estimation. While these methods can be highly accurate in specific scenarios, they often demand substantial computational resources and can become unstable in dynamic environments.
One the other hand, feature-based methods, exemplified by the ORB-SLAM series [4, 5], focus on identifying and utilizing key image features, i.e. keypoints. These methods extract keypoints with distinctive properties, compute descriptors, and perform feature matching, outlier removal, and relative motion tracking. Their back-end processes include loop-closure detection and global optimization, crucial for refining and maintaining the system’s state estimates over time. Feature-based approaches typically excel in computational efficiency and robustness across diverse scenarios. However, they may struggle in environments with significant lighting changes or insufficient texture, where traditional handcrafted keypoints might not be detected or matched reliably.
In recent years, deep learning has demonstrated remarkable capabilities in computer vision tasks, such as image classification and object detection. Its integration into Visual Odometry (VO) and SLAM has garnered considerable interest. End-to-end methods use deep neural networks (DNNs) to directly infer pose estimates from images [6, 7]. However, these methods often lack interpretability and may have difficulties generalizing across different scenarios due to discrepancies between training and testing data distributions. Learned-feature-based SLAM methods [8, 9, 10] replace traditional feature detection with deep learning-based visual feature extraction, which is then integrated into the SLAM back-end. These methods leverage both deep learning and geometric models to enhance SLAM performance.
Robust keypoint tracking and matching are vital for the reliability and effectiveness of visual SLAM systems in various applications. Although previous research has explored data-driven feature extraction in SLAM systems, there is still a need for more effective feature extraction networks tailored to the dynamic conditions encountered in SLAM applications. Current neural networks for feature extraction, such as SuperPoint [11], are typically trained on static object detection datasets, which may limit their performance in dynamic, real-world environments with varying lighting and movement. Additionally, achieving reliable matching for learned features and efficient loop-closing remains a significant challenge. Unlike ORB features, which encapsulate explicit low-level image information, learned features often produce high-dimensional descriptor embeddings, complicating the process of establishing matches and filtering outliers. Moreover, enabling loop-closure detection using learned features poses further difficulties. Learning-based Bag-of-Words (BoW) methods for loop-closure detection involve complex training processes, often requiring extensive offline training and substantial memory storage, and they struggle with generalization to new environments.
To address the aforementioned challenges, we propose DK-SLAM, a novel Deep Keypoint based SLAM system, that incorporates deep keypoint extraction through meta-learning, coarse-to-fine feature tracking, and an efficient online learning-based loop-closing. Our DK-SLAM employs an enhanced Model-Agnostic Meta-Learning (MAML) strategy to extract robust deep local features, improving the system’s ability to perform feature extraction in diverse and dynamic environments. This meta learning strategy enables our DK-SLAM system to adapt to diverse conditions without the need for additional training during deployment, thereby enhancing its applicability in real-world scenarios. Additionally, we introduce a novel feature matching strategy tailored for deep learned keypoints. This strategy involves a two-stage process: initial estimation of relative poses using patch photometric loss optimization, followed by refinement through 3D-2D relationships, resulting in improved matching accuracy and robustness. Furthermore, to enable more effective and efficient loop-closing, we propose an online learning-based Bag-of-Words (BoW) model. Our model compresses learned features into binary descriptors and utilizes images from previous timesteps within an online learning framework. It constructs a dynamic tree structure for bag-of-words, facilitating loop-closure detection in new environments.
Experimental evaluation conducted on public car-driving and drone datasets, demonstrates the superior performance of our proposed DK-SLAM compared to representative traditional visual SLAM such as ORB-SLAM3 and LDSO, as well as learning-based systems such as LIFT-SLAM. Notably, compared to the representative monocular ORB-SLAM3, DK-SLAM achieves a translation accuracy improvement of approximately 17.7% and a rotation accuracy enhancement of 24.2% on the KITTI dataset, and it surpasses ORB-SLAM3 by approximately 34.2% in translation accuracy on the EuRoC dataset.
@@
  • We propose DK-SLAM, a novel monocular visual SLAM system with deep keypoint meta learning. Our deep feature extractor, trained using a MAML strategy, enhances adaptability to diverse scenes.
  • We develop a coarse-to-fine two-stage keypoint matching strategy that estimates relative poses through patch photometric loss optimization and refines them based on the 3D-2D relationship, resulting in improved accuracy.
  • We introduce an online learning-based Bag-of-Words (BoW) model that effectively and efficiently utilizes binarized deep keypoints. This model dynamically fine-tunes itself to ensure accurate loop detection in the long-term operation.
The reminder of the paper is organized as follows: Section II contains a survey of related work; Section III introduces our proposed deep keypoint based monocular SLAM framework; Section IV evaluates DK-SLAM in both car-driving and drone scenarios, and conducts extensive ablation studies; Section V finally draws conclusions.
@@ II Related Works
In this section, we provide a brief overview of some related work in learning based feature extraction, learning based visual SLAM, and loop-closure detection.
II-A Learning based Feature Extractor
@@
Mainstream handcrafted extractors, including ORB [12], SIFT [13], and Shi-Thomas [14], are employed in visual SLAMs like ORB-SLAM2 [4] and VINS-Mono [15]. However, these methods relying on gradient information face challenges in dynamic lighting and low-texture environments. To address these issues, researchers explored deep feature extractors. Superpoint [11] employs a shared encoder for high-level information and distinct decoders for detectors and descriptors, using synthesized labels for self-supervised training. D2-Net [16] eliminates the need for a separate keypoint decoder, ensuring robustness. To enhance repeatability detection, [17] combines multi-scale gradients with depth feature maps. Additionally, L2-Net [18] introduces a local matching loss based on the L2 distance, improving feature matching performance. Experimental results demonstrate the effective feature matching performance of the learned local features on public datasets. Despite their success, the mentioned deep extractors exhibit inherent flaws in the face of diverse scenes, leading to potential performance degradation.
Figure 1: An overview of our proposed DK-SLAM framework with deep keypoint meta learning, two-stage coarse-to-fine keypoint tracking and online learning based binary BoW for loop-closing.
II-B Learning based Visual SLAM
@@
Recent research explores novel SLAM methods that exploit the advantages of both deep learning and multi-view geometry. GCNv2-SLAM integrates Graph Convolutional Networks (GCN) [19] into SLAM, culminating in a comprehensive system with offline Bag-of-Words (BoW) training. In another study [20], knowledge distillation from HF-Net [21] enhances feature detection within a compact model. Additionally, [22] employs a self-supervised thermal optical flow tracking network, while [23] enhances Superpoint [11] with edge and multi-scale information. In [9], a LIFT [24] extractor serves as the SLAM front-end, implementing an adaptive feature matching strategy tailored to different datasets. In contrast, DK-SLAM adopts a meta-learning based training strategy, obviating the need for separate hyperparameter designs for distinct datasets. Moreover, [25] proposes a binary learning feature-based VSLAM that uses an adaptive motion model to provide more accurate initial pose estimates for tracking threads, achieving commendable localization performance on public datasets. To address the challenges posed by dynamic objects in visual SLAM, [26] introduces a semantic segmentation network that identifies dynamic objects in the scene. The system then removes feature points associated with these objects during the front-end processing, leading to state-of-the-art performance in environments with multiple moving objects. Additionally, [27] exploits geometric information from objects to enhance VSLAM performance, and [28] applies learning-based semantic segmentation to correct scale in monocular SLAM mapping and localization. Furthermore, [29] incorporates learned depth labels into monocular visual SLAM, transforming it into a virtual stereo SLAM to overcome scale ambiguity. These depth labels are generated from unsupervised depth estimation networks, enabling the system to achieve more accurate mapping and localization.
II-C Loop Closure Detection
Loop-closure detection is a crucial component of visual SLAM, tasked with identifying loop nodes and refining initial pose estimates [30]. Traditional Bag-of-Words (BoW) models like DBow2, DBow3, and FBow [31] utilize the binary descriptor BRIEF [32] and store descriptors in a k-d tree structure based on Hamming distance. Despite their utility, these conventional BoW methods require extensive offline training data, which can limit their adaptability. To address this, iBow [33] was introduced, offering online BoW training directly from testing datasets. This approach provides robust transfer performance across various scenarios, enhancing the system’s adaptability. Another approach to loop-closure detection, similar to NetVLAD [34], as used in [20, 10], depends heavily on the quality and relevance of the training set. Significant differences between the training and testing datasets can lead to performance degradation. In contrast, [35] utilizes deep learning to handle loop closure under varying viewpoints and lighting conditions. Instead of relying on the entire image, this approach focuses on landmarks to determine loop closure, enhancing its robustness and accuracy in diverse environments.
@@ III Deep Keypoint based Monocular SLAM
An overview of our proposed DK-SLAM system is illustrated in Figure 1. In the front-end, we utilize a neural network-based feature extractor to detect and describe keypoints across multi-scale image pyramids. To enhance generalization, the local feature extractor undergoes Model-Agnostic Meta-Learning (MAML) during training. To ensure a balanced distribution of keypoints across images of varying scales, we implement an averaging distribution strategy. Subsequently, to enhance the tracking and matching performance of the learned visual keypoints, we incorporate low-level information from the image. Initially, we construct local photometric constraints on the pixels surrounding keypoints to estimate the initial relative pose. Furthermore, we estimate the matching range of keypoints in adjacent frames based on this initial relative pose. Once the front-end completes keypoint tracking and pose estimation, keyframes are selected based on a predefined interval time scale. These selected keyframes are imported into the keyframe database, with the most recent keyframe being integrated into the back-end loop closure module. Memory consumption is optimized by converting floating keyframe descriptors into binary hash codes. Lastly, we introduce an online Bag-of-Words (BoW) module to establish a loop closure mechanism, facilitating closed-loop detection and global map optimization.
III-A Deep Keypoint Meta Learning
@@ III-A1 Feature Extractor Network
Inspired by SuperPoint [11], our deep keypoint extractor also leverages VGG16 [36] as the backbone architecture. Unlike SuperPoint, we introduce Batch Normalization layers after each Convolutional Neural Network (CNN) to encourage training convergence. After this backbone, there are decoders for keypoint detector and descriptor, each consisting of multi-layer CNNs. The first module, i.e. the detector decoder, produces a probability mask matching the input image size. We posit that keypoints are situated at positions where the mask value surpasses a certain threshold. The second module, i.e. the descriptor decoder, predicts feature descriptors for keypoints, which are then normalized to a dimension of 256 using the L2 norm.
III-A2 Self-Supervised Keypoint Learning
Our network undergoes self-supervised training, employing a multi-step process. Initially, we obtain the homography-warped images from current training images. To train keypoint detector, we leverage the pre-trained MagicPoint network [11] to acquire keypoint pseudo labels for both the original and warped images. In both images, keypoint predictions and pseudo labels contribute to the formation of detector loss, denoted as ${L_{p}}$ and ${L_{wp}}$, with CrossEntropy serving as the loss metric. ${L_{p}}$ represents the detector loss of the original image, while $L_{wp}$ represents the detector loss of the warped image. Moving on to descriptor training, we utilize the sparse descriptor loss ${L_{d}}$ proposed in [37]. This involves obtaining both the warped image and pixel-wise correspondences. Descriptor loss functions are then constructed solely for pseudo-labeled keypoints, employing triplet loss. The sparse descriptor loss is composed of two components: homography-matched descriptor loss $L_{dm}$ and non-matched descriptor loss $L_{dn}$. The homography-matched descriptor loss, represented by Equation 1, involves a subscript $i$ ranging between $(0,1,...,N)$, where $(\mathbf{u}_{i},\hat{\mathbf{u}_{i}})$ denotes the $N$ matched keypoint position pairs.
$L_{dm}(\mathbf{u}_{i})={\left\|\mathbf{d}_{I}(\mathbf{u}_{i})-\mathbf{d}_{wI}(\hat{\mathbf{u}_{i}})\right\|}^{2}$ (1)
As depicted in Equation 2, we use $M$ pairs of descriptors from the vicinity of matched descriptors to formulate a non-matched descriptor loss. The subscript $j$ ranges from 0 to $M$, where $(\mathbf{u}_{i},\mathbf{u^{\prime}}_{j})$ represents unmatched keypoint position pairs.
$L_{dn}(\mathbf{u}_{i})=\frac{1}{M}\sum_{j=1}^{M}(\left\|\mathbf{d}_{I}(\mathbf{u}_{i})-\mathbf{d}_{wI}(\mathbf{u^{\prime}}_{j})\right\|^{2})$ (2)
@@ III-A2 Self-Supervised Keypoint Learning
III-A3 MAML-Based Visual Keypoint Meta Learning
@@
Learning-based visual keypoint (local feature) extractors face poor generalization due to variances in visual features across scenes. Significant feature disparity between training and unseen datasets can lead to catastrophic forgetting. To address this, we leverage insights from Model-Agnostic Meta-Learning (MAML), adapting network parameters through meta-training, which enhances generalization to new scenarios without requiring additional training during deployment. The MAML-based meta-training consists of inner loop training and outer loop training. The training set is partitioned into a support set $D_{s}$ and a query set $D_{q}$, where $D_{s}$ is involved in inner loop training, and $D_{q}$ participates in outer loop training.
@@
In the MAML training strategy, the original network parameter $\mathbf{\theta}_{a}$ is initially duplicated to $\mathbf{\theta}_{b}$. Subsequently, both parameters undergo training in both the inner and outer loops. Within the inner loop, the support dataset $D_{s}$ is utilized for iterative updates to $\mathbf{\theta}_{a}$. This involves dividing a batch of support dataset into $n$ distinct tasks $(D^{i}_{s},i=0...n-1)$, performing $m$ parameter updates on each task, resulting in the updated parameter $\mathbf{\theta}^{m}_{a}$. After completing one iteration of inner loop training, a minibatch of query set data $D^{i}_{q}$ is used to update the $\mathbf{\theta}_{b}$. The detailed steps of this feature meta-training strategy using MAML are shown in the Algorithm 1 below.
Algorithm 1 The Procedure of Deep Keypoint Meta Learning in DK-SLAM.
1:Feature training dataset $D_{f}$, inner loop learning rate $\alpha_{1}$, outer loop learning rate $\alpha_{2}$.
2:Copy the initial network parameters as $\mathbf{\theta}_{a}$ and $\mathbf{\theta}_{b}$.
3:while not done do
4:     Sample a batch of support set $D_{s}$ and query set $D_{q}$ from $D_{f}$, whose batchsize are the same
5:     Sample a minibatch $D^{i}_{s}$ from $D_{s}$. Sample a minibatch $D^{i}_{q}$ from $D_{q}$. Where $i=0...n-1$. $n$ is the number of batchsize.
6:     for $D^{i}_{s}$ and $D^{i}_{q}$ in $D_{s}$, $D_{q}$ do
7:         Calculate the gradient on $\mathbf{\theta}_{a}$ using $D^{i}_{s}$ and equation(4). The gradient is $\bigtriangledown_{\mathbf{\theta}_{a}}L_{all}$.
8:         Update $\mathbf{\theta}_{a}$ with $\bigtriangledown_{\mathbf{\theta}_{a}}L_{all}$. $\mathbf{\theta}^{\prime}_{a}=\mathbf{\theta}_{a}-\alpha_{1}\bigtriangledown_{\mathbf{\theta}_{a}}L_{all}$
9:         Repeat step 6 and 7 for m iterations. Received updated m times $\mathbf{\theta}^{m}_{a}$. In our implementation, we set m as 4. We call these iterations as inner loop.
10:         Calculate the gradient on $\mathbf{\theta}^{m}_{a}$ using $D^{i}_{q}$ and equation(4). The gradient is $\bigtriangledown_{\mathbf{\theta}^{m}_{a}}L_{all}$.
11:         Update $\mathbf{\theta}_{b}$ with $\bigtriangledown_{\mathbf{\theta}^{m}_{a}}L_{all}$. $\mathbf{\theta}^{\prime}_{b}=\mathbf{\theta}_{b}-\alpha_{2}\bigtriangledown_{\mathbf{\theta}^{m}_{a}}L_{all}$ We call the updating for $\mathbf{\theta}_{b}$ as outer loop.
12:     end for
13:end while
III-A4 The Distribution Strategy of Deep Keypoints